OpenStudy (anonymous):

evaluate f(-2) and f(4). Find the zero of f. f(x)=2-1/2x

OpenStudy (apolloschariot):

Can you post the equation?

OpenStudy (anonymous):

@ApollosChariot I just did srry.

OpenStudy (apolloschariot):

Oh, okay, I didn't even see it. Basically what you do is when it says f(-2), you plug -2 into the equation for each x. And for f(4), you plug 4 into the equation for each x

OpenStudy (anonymous):

\[f \left( x \right)=2-\frac{ 1 }{ 2 }x\] how do you find the zero of f though @ApollosChariot

OpenStudy (apolloschariot):

You set the equation equal to 0 and then solve for x

OpenStudy (skullpatrol):



OpenStudy (anonymous):

now it is much more clear find the zero by setting it equal to zero and solve

OpenStudy (anonymous):

ok last thing how would I graph -4 which is the zero @satellite73

OpenStudy (anonymous):

ok lets take it from the top\[f \left( x \right)=2-\frac{ 1 }{ 2 }x\]

OpenStudy (anonymous):

put \[2-\frac{ 1 }{ 2 }x=0\] solve for \(x\)

OpenStudy (anonymous):

\[2-\frac{1}{2}x=0\\ -\frac{1}{2}x=-2\\ x=-2\times -2=4\]

OpenStudy (anonymous):

point on the graph is \((4,0)\)

OpenStudy (skullpatrol):

Why do you think the point is called a "zero" of the graph?